# A Review of Graph Neural Networks in Epidemic Modeling

### v1

@@ 3.1.3. Prediction
    Incidence Prediction. The target of incidence prediction is to provide quantitative results. In epidemic forecasting, incidences can include the number of infections or deaths in the future  (sijiraniSpatioTemporalPredictionEpidemiology2023; yu2023spatiotemporal; nguyenPredictingCOVID19Pandemic2023; croftForecastingInfectionsSpatiotemporal2023; puDynamicAdaptiveSpatio2023; qiu2023msgnna; moonRESEATRecurrentSelfAttention2023; caoMetapopulationGraphNeural2023; tangEnhancingSpatialSpread2023), influenza activity level  (liuEpidemiologyawareDeepLearning2023), Influenza-Like Illness (ILI) rates  (zhangPredictingInfluenzaPandemicawareness2023), vaccine hesitancy  (moonGraphBasedDeep2023), etc. The prediction of these incidences is important to decision-making, proactive public health planning, and the effective management of infectious diseases and other health challenges.
@@
    Trend Prediction. Different from incidence prediction task, which focuses on quantitative targets, the target of trend prediction tasks is to identify a higher-level epidemic spreading pattern. For transmission among locations  (liu2023human), prediction of infection trend can be described as an information retrieving problem and the goal is to predict the next region to be infected given a historic spreading route. For transmission among individuals or groups, the goal usually includes identifying transmission dynamics in emerging high-risk groups  (sunDeepDynaForecastPhylogeneticinformedGraph2023).
@@ 3.1.4. Projection
In epidemic forecasting, projection tasks are similar to prediction, but with an additional intention to understand epidemic outcomes. These tasks usually require models with the ability to incorporate changes during the evolving of epidemics, such as external interventions and changing of initial states. Most of the projection tasks we collected involve finding the optimal interventions or maximizing influence to achieve targets like curbing the spread of diseases. Influence maximization  (kempe2003maximizing) aims to identify a subset of nodes so that the influence spreads most effectively across the graph, and there have been several early studies in epidemiological tasks, e.g., node importance ranking  (bucur2020beyond; holme2017three).
@@
For traditional influence maximization tasks and vaccine strategy tasks  (jhun2021effective), which aim to vaccinate the optimal set of nodes to minimize epidemic damage, strategies are limited to nodes at the starting time point, which is $P_{v}(0)$. For interventions throughout the period, strategies can include applying quarantine level to nodes at each step  (fengContactTracingEpidemic2023), which denotes $P_{v}(t)$ or restricting mobility on edges  (meirom2021controlling; song2020reinforced), which denotes $P_{\mathcal{E}}(t)$.
@@
    For inputs of all models, they all consist mainly of two parts: node features and the graph structure. Based on the temporality of nodes, we can further categorize these work into spatial-only tasks, temporal-only   (zhangPredictingInfluenzaPandemicawareness2023), and spatial-temporal tasks. In addition, based on the temporality and learnability of graph structure, we can also use static or dynamic features to distinguish these works. A detailed introduction is presented below in Section 3.2.
@@
    In terms of model outputs, there are also three categories to summarize these works: scalar, graph, and action sequence. Scalar outputs are usually used in prediction tasks which provide indicators of the epidemic like infected cases. There are also some works that focus on epidemic graph construction and the outputs of their designed models are graphs  (shan2023novel; wang2023wdcip). Finally, the projection tasks we collected usually adopt Reinforcement Learning (RL), which outputs the actions taken at each time step, forming a consecutive action sequence  (jhun2021effective; fengContactTracingEpidemic2023; meirom2021controlling; song2020reinforced).
@@ 3.2. Data Sources
The investigated works in this survey cover a wide range of datasets from different parts of the world, as shown in Figure  2. However, the majority of research focuses only on COVID-19 while only a few study Influenza-Like Illnesses (ILI) or bacteria  (gouarebDetectionPatientsRisk2023). We categorize these datasets based on their sources:
@@
    Online Search and Social Media. Epidemic information can also be acquired through social media and online search records. The massive search of disease-related questions in a region can indicate potential outbreaks (lin2023grapha), which can then be utilized as node features.
@@
    Sensors. Multi-modal data can be acquired through sensors like cameras, satellites, radios, etc. These data can also help epidemic tasks like exposure risk prediction using images  (hanDevilLandscapesInferring2023). Unlike conventional data sources, sensor data often requires additional preprocessing using specialized models, such as encoding images with techniques like ResNet (he2016deep), before integration as node features.
@@
    Simulated Data. Besides real-world data, some research  (tang2023enhancing; sunDeepDynaForecastPhylogeneticinformedGraph2023; meznar2021prediction) also utilized simulated data for model training and testing. These data often require simulation models like TimeGEO  (jiang2016timegeo), Independent Contagion Model (ICM)  (murphy2021deep), and also SIR models to generate temporal graphs.
@@ 3.3. Graph Construction
Static node features typically refer to characteristics that do not change with time. The shape of static features can be represented as $\mathbb{R}^{N\times h}$, where $h$ refers to the number of different features. Besides tasks involving time series, most GNN tasks are using static features. For example, in a contact graph in which individuals are modeled as nodes and contact information represents edges, personal characteristics like infection order, gender, age, and symptom-related information can be used as static features during training and prediction  (songCOVID19InfectionInference2023). However, tasks involving time series can also use static features as additional information.
@@
Contrary to static features, dynamic features represent characteristics that change through time. This type of data is commonly seen in time-series forecasting tasks and the models usually require inputs at each time point. Therefore, the shape of the dynamic features can be represented as $\mathbb{R}^{N\times T\times h}$, where $T$ refers to the number of time points given. As an example, the number of daily confirmed cases in a region can be seen as dynamic features  (xieEpiGNNExploringSpatial2022). Although most models take in a single slice of dynamic features at each time point, some models use the entire dynamic features across time $T$ in a single input  (ruInferringPatientZero2023).
@@
The construction of a static graph structure typically entails the use of a predefined approach to generate the graph from available data. Once the graph is established, its structure remains unchanged throughout the training iterations or over different time points. For instance, in tasks involving multiple regions, the geographical adjacency $A$, is often employed to connect different regions, which are represented as nodes in the graph $G$  (pu2023dynamic; yu2023spatiotemporal). The distance between regions or other features can be considered as the edge weights. Another strategy focuses on exploring human mobility or transitions, e.g., linking nodes through nearest neighbors in the case of COVID-19 transmission. This method takes into consideration the distribution of the population and individual movements between various locations  (liu2023human; wang2023wdcip). In research where the node represents an individual, connections between two nodes often utilize contact information, e.g., identifying contacts at risk of spreading the disease as links between individuals  (gouareb2023detectiona; tomy2022estimating).
@@
Determining the structure of a dynamic graph commonly involves one of two methodologies. One approach is the modification of adjacency relations over time or throughout the virus propagation process. For example,  (meirom2021controlling) utilize $\mathcal{E}(t)=\{e_{uv}(t)\}$ to represent the set of edges at time step $t$, which connect individuals based on transmission probability. Another strategy entails the learning of adaptive edges or edge weights during the training phase. Given the dynamic nature of disease transmission, which evolves at each time step, traditional geographical adjacency matrices fall short of accurately representing true connectivity. Recent studies  (wangCausalGNNCausalBasedGraph2022; deng2020colagnn; caoMepoGNNMetapopulationEpidemic2022) have aimed for models to learn an adaptive relationship between nodes. This typically involves initially generating node features via a neural network, followed by the computation of an attention matrix to depict dynamic connectivity, often expressed as ${\bf A}_{t}=a_{i,j}\in\mathbb{R}^{N\times N}$, where $a_{i,j}$ indicates the influence of node $v_{j}$ on node $v_{i}$.
@@ 4.1.1. Spatial Dynamics Modeling
One advantage of GNNs, e.g., GCN or GAT, is their ability to capture spatial relationships through various aggregation processes, which can analyze and capture the spatial dimensions of disease propagation. Numerous studies represent the inherent structure of geographical data as graph data, denoted as ${\bf A}$, where nodes depict regions (e.g., cities, neighborhoods, or countries), and edges describe connections between these regions (e.g., roads, flights, or potential vectors for disease transmission). Subsequently, GNNs are applied to the graph data to uncover complex relationships and dependencies at the regional level, facilitating predictions regarding disease spread across different areas (song2023covid19; pudynamic; nguyen2023predicting; gouareb2023detectiona; lagatta2021epidemiological).
In the context of GNN modeling, the significance of edge weights is paramount, as they encapsulate the intensity and nature of interactions. Within epidemiological studies, these weights are often derived from the mobility or social connectedness between regions  (lin2023grapha; panagopoulos2021transfer). For instance, studies such as  (song2020reinforced; cao2023metapopulation) utilize Origin-Destination (OD) flows to quantify inter-regional mobility, thereby dynamically capturing the intensity of transmission. To further enhance the spatial context of each node within the graph, some research advocates for the implementation of positional encoding techniques  (wang2022equivariant; bograph). These techniques are designed to augment the nodes’ spatial awareness. For example, Liu et al. (liu2023human) introduced a unique encoding for each location, denoted as $PE(k)$, with even and odd elements represented by $\sin(k/10000^{i/L})$ and $\cos(k/10000^{i-1/L})$ respectively, where $L$ denotes the dimension of the encoding.
@@
Adaptive Structure Learning. Although GNNs possess the inherent capability to learn the spatial characteristics of disease dissemination, there are occasions when adjacency relationship information is not available in the real world, often due to data scarcity. To overcome this challenge, several studies highlight the importance of learning an adaptive structure throughout the training process (zheng2021hierst; wang2022adaptively). For instance, Wang et al. (wang2022adaptively) introduced a graph structure learning module, denoted as $f_{\theta}$. This module is designed to calculate node similarities, thereby representing spatial relationships as follows:
@@
where $\mathbf{X}_{1}$ and $\mathbf{X}_{2}$ are randomly initialized, learnable node embeddings, while $\alpha$ represents a hyper-parameter. Shan et al. (shan2023novel) employed a method to estimate the graph Laplacian from COVID-19 data through convex optimization of derived eigenvectors. This approach aims to identify dynamic patterns of pandemic spread among countries by analyzing their structural relationships. Additionally, inspired by recent advancements in attention-based mechanisms (vaswani2023attention; turner2024introduction; soydaner2022attention), a considerable portion of research suggests the use of an attention matrix to illustrate the relationships between nodes (deng2020colagnn; wangCausalGNNCausalBasedGraph2022; cui2021unobservables). Notably, Cola-GNN (deng2020colagnn) pioneers the application of additive attention in learning the adaptive structure, which is defined as follows:
@@
Multi-Scale Modeling. Previous approaches typically operate at a singular level, overlooking the multifaceted nature of real-world epidemiological data, which encompasses multiple scales such as country, state, and community levels. For epidemiological tasks, multi-scale modeling is imperative for capturing the dynamics of disease spread across these varied levels, from individual behaviors to global dissemination, ensuring a more comprehensive analysis. HierST (zheng2021hierst) leverages multi-scale modeling to effectively capture the spread of COVID-19 across different administrative levels by constructing a unified graph, which encapsulates the spatial correlation dynamics both within and between these levels. To further address both local interactions and long-range dependencies, MSGNN (qiu2023msgnna) is meticulously designed to integrate influences from both immediate and broader regions on disease transmission, enhancing its effectiveness in cross-scale epidemiological dynamics.
@@ 4.1.2. Temporal Dynamics Modeling
The temporal dynamics in epidemiological models are pivotal for capturing the evolution of disease spread, reflecting changes in infection rates, recovery rates, and other critical parameters over time. These models typically conceptualize graphs as spatio-temporal networks, underscoring the significance of temporal data in comprehending disease dynamics, and forecasting future trends (zhang2023predicting; sijirani2023spatiotemporala). A particular strand of research utilizes RNN-based models (e.g., LSTM or GRU), as mechanisms to extract node features. These features are then incorporated into the graph convolution process (CoLA_TNNLS21; zheng2021hierst). A simple way (kapoor2020examining) to achieve this by executing the concat operator:
@@
where h is simply the output of an MLP (Multilayer Perceptron) over the node temporal features x at time $t$ reaching back $d$ days. Another surge of approaches first executes graph spatial convolution in each time step separately, and then leverages all outputs of the GNNs as the input of the temporal module and utilizes them for final downstream tasks like the prediction (gao2021stan; sijirani2023spatiotemporala; moon2023graph; duarte2023time). STEP (sijirani2023spatiotemporala) execute the multi-layers graph convolution operation to get all node embedding h, and then leverage the GRU to get the final output:
@@
In contrast to the initial two methodologies, numerous studies achieve their final output by iteratively layering GNN and temporal models (sesti2021integrating). Some work (guo2021research; sha2021source) advocate for the employment of Spatio-Temporal Graph Neural Networks (STGNNs) (hu2022graph; zhou2021graph; wu2021comprehensive) to extract insights from multivariate spatiotemporal epidemic graphs. An STGNN integrates many ST-Conv blocks, which comprise a spatial layer flanked by two temporal layers. Each temporal layer features a 1-D CNN operating along the time axis, followed by a Gated Linear Unit (GLU), to delineate the temporal dynamics. The spatial layer, on the other hand, utilizes a GCN based on the Chebyshev polynomials approximation (Cheb_Neurips16; Cheb2_Neurips22) for spatial analysis. To further refine the understanding of spatial dynamics during disease evolution, RESEAT  (moon2023reseat) proposes the continuous maintenance and adaptive updating of an attention matrix. This process aims to capture regional interrelationships throughout the entirety of the input data period:
@@
The $\textbf{A}^{t}_{i,j}\in\mathbb{R}$ denotes the attention weight between regions $i$ and $j$ at time step $t$, and $Att_{i}$ is employed as the final feature for the node $v_{i}$. Through this mechanism, RESEAT adeptly captures not only temporal patterns but also the dynamically evolving regional interrelationships. To integrate explicit observations with implicit factors over time, Cui et al. (cui2021unobservables) introduced a new case prediction methodology within an encoder-decoder framework. They contend that relying solely on observed case data, which can be inaccurate, may impair prediction performance. Accordingly, their proposed decoder is designed to incorporate inputs of new cases and deaths, thereby dynamically reflecting temporal changes.

### v2

@@ 3.1.3. Prediction
    Incidence Prediction. The target of incidence prediction is to provide quantitative results. In epidemic forecasting, incidences can include the number of infections or deaths in the future  (sijiraniSpatioTemporalPredictionEpidemiology2023, 79, 80, 81, 82, 66, 67, 83, 84, 85), influenza activity level  (liuEpidemiologyawareDeepLearning2023, 86), Influenza-Like Illness (ILI) rates  (zhangPredictingInfluenzaPandemicawareness2023, 87), vaccine hesitancy  (moonGraphBasedDeep2023, 88), etc. The prediction of these incidences is important to decision-making, proactive public health planning, and the effective management of infectious diseases and other health challenges.
@@
    Trend Prediction. Different from incidence prediction task, which focuses on quantitative targets, the target of trend prediction tasks is to identify a higher-level epidemic spreading pattern. For transmission among locations  (liu2023human, 61), prediction of infection trend can be described as an information retrieving problem and the goal is to predict the next region to be infected given a historic spreading route. For transmission among individuals or groups, the goal usually includes identifying transmission dynamics in emerging high-risk groups  (sunDeepDynaForecastPhylogeneticinformedGraph2023, 89).
@@ 3.1.4. Projection
In epidemic forecasting, projection tasks are similar to prediction, but with an additional intention to understand epidemic outcomes. These tasks usually require models with the ability to incorporate changes during the evolving of epidemics, such as external interventions and changing of initial states. Most of the projection tasks we collected involve finding the optimal interventions or maximizing influence to achieve targets like curbing the spread of diseases. Influence maximization  (kempe2003maximizing, 90) aims to identify a subset of nodes so that the influence spreads most effectively across the graph, and there have been several early studies in epidemiological tasks, e.g., node importance ranking  (bucur2020beyond, 91, 92).
@@
For traditional influence maximization tasks and vaccine strategy tasks  (jhun2021effective, 59), which aim to vaccinate the optimal set of nodes to minimize epidemic damage, strategies are limited to nodes at the starting time point, which is $P_{v}(0)$. For interventions throughout the period, strategies can include applying quarantine level to nodes at each step  (fengContactTracingEpidemic2023, 60), which denotes $P_{v}(t)$ or restricting mobility on edges  (meirom2021controlling, 93, 30), which denotes $P_{\mathcal{E}}(t)$.
@@
    For inputs of all models, they all consist mainly of two parts: node features and the graph structure. Based on the temporality of nodes, we can further categorize these work into spatial-only tasks, temporal-only   (zhangPredictingInfluenzaPandemicawareness2023, 87), and spatial-temporal tasks. In addition, based on the temporality and learnability of graph structure, we can also use static or dynamic features to distinguish these works. A detailed introduction is presented below in Section 3.2.
@@
    In terms of model outputs, there are also three categories to summarize these works: scalar, graph, and action sequence. Scalar outputs are usually used in prediction tasks which provide indicators of the epidemic like infected cases. There are also some works that focus on epidemic graph construction and the outputs of their designed models are graphs  (shan2023novel, 94, 58). Finally, the projection tasks we collected usually adopt Reinforcement Learning (RL), which outputs the actions taken at each time step, forming a consecutive action sequence  (jhun2021effective, 59, 60, 93, 30).
@@ 3.2. Data Sources
The investigated works in this survey cover a wide range of datasets from different parts of the world, as shown in Figure  2. However, the majority of research focuses only on COVID-19 while only a few study Influenza-Like Illnesses (ILI) or bacteria  (gouarebDetectionPatientsRisk2023, 78). We categorize these datasets based on their sources:
@@
    Online Search and Social Media. Epidemic information can also be acquired through social media and online search records. The massive search of disease-related questions in a region can indicate potential outbreaks (lin2023grapha, 64), which can then be utilized as node features.
@@
    Sensors. Multi-modal data can be acquired through sensors like cameras, satellites, radios, etc. These data can also help epidemic tasks like exposure risk prediction using images  (hanDevilLandscapesInferring2023, 50). Unlike conventional data sources, sensor data often requires additional preprocessing using specialized models, such as encoding images with techniques like ResNet (he2016deep, 95), before integration as node features.
@@
    Simulated Data. Besides real-world data, some research  (tang2023enhancing, 63, 89, 96) also utilized simulated data for model training and testing. These data often require simulation models like TimeGEO  (jiang2016timegeo, 97), Independent Contagion Model (ICM)  (murphy2021deep, 98), and also SIR models to generate temporal graphs.
@@ 3.3. Graph Construction
Static node features typically refer to characteristics that do not change with time. The shape of static features can be represented as $\mathbb{R}^{N\times h}$, where $h$ refers to the number of different features. Besides tasks involving time series, most GNN tasks are using static features. For example, in a contact graph in which individuals are modeled as nodes and contact information represents edges, personal characteristics like infection order, gender, age, and symptom-related information can be used as static features during training and prediction  (songCOVID19InfectionInference2023, 77). However, tasks involving time series can also use static features as additional information.
@@
Contrary to static features, dynamic features represent characteristics that change through time. This type of data is commonly seen in time-series forecasting tasks and the models usually require inputs at each time point. Therefore, the shape of the dynamic features can be represented as $\mathbb{R}^{N\times T\times h}$, where $T$ refers to the number of time points given. As an example, the number of daily confirmed cases in a region can be seen as dynamic features  (xieEpiGNNExploringSpatial2022, 75). Although most models take in a single slice of dynamic features at each time point, some models use the entire dynamic features across time $T$ in a single input  (ruInferringPatientZero2023, 76).
@@
The construction of a static graph structure typically entails the use of a predefined approach to generate the graph from available data. Once the graph is established, its structure remains unchanged throughout the training iterations or over different time points. For instance, in tasks involving multiple regions, the geographical adjacency $A$, is often employed to connect different regions, which are represented as nodes in the graph $G$  (pu2023dynamic, 99, 80). The distance between regions or other features can be considered as the edge weights. Another strategy focuses on exploring human mobility or transitions, e.g., linking nodes through nearest neighbors in the case of COVID-19 transmission. This method takes into consideration the distribution of the population and individual movements between various locations  (liu2023human, 61, 58). In research where the node represents an individual, connections between two nodes often utilize contact information, e.g., identifying contacts at risk of spreading the disease as links between individuals  (gouareb2023detectiona, 100, 101).
@@
Determining the structure of a dynamic graph commonly involves one of two methodologies. One approach is the modification of adjacency relations over time or throughout the virus propagation process. For example,  (meirom2021controlling, 93) utilize $\mathcal{E}(t)=\{e_{uv}(t)\}$ to represent the set of edges at time step $t$, which connect individuals based on transmission probability. Another strategy entails the learning of adaptive edges or edge weights during the training phase. Given the dynamic nature of disease transmission, which evolves at each time step, traditional geographical adjacency matrices fall short of accurately representing true connectivity. Recent studies  (wangCausalGNNCausalBasedGraph2022, 55, 23, 26) have aimed for models to learn an adaptive relationship between nodes. This typically involves initially generating node features via a neural network, followed by the computation of an attention matrix to depict dynamic connectivity, often expressed as ${\bf A}_{t}=a_{i,j}\in\mathbb{R}^{N\times N}$, where $a_{i,j}$ indicates the influence of node $v_{j}$ on node $v_{i}$.
@@ 4.1.1. Spatial Dynamics Modeling
One advantage of GNNs, e.g., GCN or GAT, is their ability to capture spatial relationships through various aggregation processes, which can analyze and capture the spatial dimensions of disease propagation. Numerous studies represent the inherent structure of geographical data as graph data, denoted as ${\bf A}$, where nodes depict regions (e.g., cities, neighborhoods, or countries), and edges describe connections between these regions (e.g., roads, flights, or potential vectors for disease transmission). Subsequently, GNNs are applied to the graph data to uncover complex relationships and dependencies at the regional level, facilitating predictions regarding disease spread across different areas (song2023covid19, 102, 103, 65, 100, 104).
In the context of GNN modeling, the significance of edge weights is paramount, as they encapsulate the intensity and nature of interactions. Within epidemiological studies, these weights are often derived from the mobility or social connectedness between regions  (lin2023grapha, 64, 105). For instance, studies such as  (song2020reinforced, 30, 70) utilize Origin-Destination (OD) flows to quantify inter-regional mobility, thereby dynamically capturing the intensity of transmission. To further enhance the spatial context of each node within the graph, some research advocates for the implementation of positional encoding techniques  (wang2022equivariant, 106, 107). These techniques are designed to augment the nodes’ spatial awareness. For example, Liu et al. (liu2023human, 61) introduced a unique encoding for each location, denoted as $PE(k)$, with even and odd elements represented by $\sin(k/10000^{i/L})$ and $\cos(k/10000^{i-1/L})$ respectively, where $L$ denotes the dimension of the encoding.
@@
Adaptive Structure Learning. Although GNNs possess the inherent capability to learn the spatial characteristics of disease dissemination, there are occasions when adjacency relationship information is not available in the real world, often due to data scarcity. To overcome this challenge, several studies highlight the importance of learning an adaptive structure throughout the training process (zheng2021hierst, 72, 71). For instance, Wang et al. (wang2022adaptively, 71) introduced a graph structure learning module, denoted as $f_{\theta}$. This module is designed to calculate node similarities, thereby representing spatial relationships as follows:
@@
where $\mathbf{X}_{1}$ and $\mathbf{X}_{2}$ are randomly initialized, learnable node embeddings, while $\alpha$ represents a hyper-parameter. Shan et al. (shan2023novel, 94) employed a method to estimate the graph Laplacian from COVID-19 data through convex optimization of derived eigenvectors. This approach aims to identify dynamic patterns of pandemic spread among countries by analyzing their structural relationships. Additionally, inspired by recent advancements in attention-based mechanisms (vaswani2023attention, 108, 109, 110), a considerable portion of research suggests the use of an attention matrix to illustrate the relationships between nodes (deng2020colagnn, 23, 55, 111). Notably, Cola-GNN (deng2020colagnn, 23) pioneers the application of additive attention in learning the adaptive structure, which is defined as follows:
@@
Multi-Scale Modeling. Previous approaches typically operate at a singular level, overlooking the multifaceted nature of real-world epidemiological data, which encompasses multiple scales such as country, state, and community levels. For epidemiological tasks, multi-scale modeling is imperative for capturing the dynamics of disease spread across these varied levels, from individual behaviors to global dissemination, ensuring a more comprehensive analysis. HierST (zheng2021hierst, 72) leverages multi-scale modeling to effectively capture the spread of COVID-19 across different administrative levels by constructing a unified graph, which encapsulates the spatial correlation dynamics both within and between these levels. To further address both local interactions and long-range dependencies, MSGNN (qiu2023msgnna, 67) is meticulously designed to integrate influences from both immediate and broader regions on disease transmission, enhancing its effectiveness in cross-scale epidemiological dynamics.
@@ 4.1.2. Temporal Dynamics Modeling
The temporal dynamics in epidemiological models are pivotal for capturing the evolution of disease spread, reflecting changes in infection rates, recovery rates, and other critical parameters over time. These models typically conceptualize graphs as spatio-temporal networks, underscoring the significance of temporal data in comprehending disease dynamics, and forecasting future trends (zhang2023predicting, 62, 112). A particular strand of research utilizes RNN-based models (e.g., LSTM or GRU), as mechanisms to extract node features. These features are then incorporated into the graph convolution process (CoLA_TNNLS21, 113, 72). A simple way (kapoor2020examining, 114) to achieve this by executing the concat operator:
@@
where h is simply the output of an MLP (Multilayer Perceptron) over the node temporal features x at time $t$ reaching back $d$ days. Another surge of approaches first executes graph spatial convolution in each time step separately, and then leverages all outputs of the GNNs as the input of the temporal module and utilizes them for final downstream tasks like the prediction (gao2021stan, 25, 112, 115, 116). STEP (sijirani2023spatiotemporala, 112) execute the multi-layers graph convolution operation to get all node embedding h, and then leverage the GRU to get the final output:
@@
In contrast to the initial two methodologies, numerous studies achieve their final output by iteratively layering GNN and temporal models (sesti2021integrating, 117). Some work (guo2021research, 118, 57) advocate for the employment of Spatio-Temporal Graph Neural Networks (STGNNs) (hu2022graph, 119, 120, 121) to extract insights from multivariate spatiotemporal epidemic graphs. An STGNN integrates many ST-Conv blocks, which comprise a spatial layer flanked by two temporal layers. Each temporal layer features a 1-D CNN operating along the time axis, followed by a Gated Linear Unit (GLU), to delineate the temporal dynamics. The spatial layer, on the other hand, utilizes a GCN based on the Chebyshev polynomials approximation (Cheb_Neurips16, 122, 123) for spatial analysis. To further refine the understanding of spatial dynamics during disease evolution, RESEAT  (moon2023reseat, 73) proposes the continuous maintenance and adaptive updating of an attention matrix. This process aims to capture regional interrelationships throughout the entirety of the input data period:
@@
The $\textbf{A}^{t}_{i,j}\in\mathbb{R}$ denotes the attention weight between regions $i$ and $j$ at time step $t$, and $Att_{i}$ is employed as the final feature for the node $v_{i}$. Through this mechanism, RESEAT adeptly captures not only temporal patterns but also the dynamically evolving regional interrelationships. To integrate explicit observations with implicit factors over time, Cui et al. (cui2021unobservables, 111) introduced a new case prediction methodology within an encoder-decoder framework. They contend that relying solely on observed case data, which can be inaccurate, may impair prediction performance. Accordingly, their proposed decoder is designed to incorporate inputs of new cases and deaths, thereby dynamically reflecting temporal changes.